Grade 10Mechanics

A physical quantity is represented by X = Ma Lb Tc . If the percentage errors in measurements of M , L and T are alpha % , beta % and gamma % respectively , then what is the total percentage error on the measurement of physical quantity ?

Aiden Mathew Denny
8 Years agoGrade 10

2 Answers

Venkat
ApprovedApproved Tutor Answer8 Years ago
Given      X  = MaLbTc
 
Percentage error in  X  =  (∆X/X)×100
=   a(∆M/M)×100   + b(∆L/L)×100 +  c(∆T/T)×100
 
But   (∆M/M)×100  =  α;  (∆L/L)×100 =  β; (∆T/T)×100  =  γ
= ( aα + bβ  +  cγ )%
ankit singh
5 Years ago

Answer

Given, X=MηLθTϕ

Differentiating, ΔX=(ηMη1ΔM)LθTϕ(θLθ1ΔL)MηTϕ(ϕTϕ1ΔT)MηLϕ
So, XΔX=ηMΔMθLΔLϕTΔT 
As the error should be both positive and negative , so we can take mod
The % error in X =XΔX×100
                            =[ηMΔM+θLΔL+ϕTΔT]×100   
                            =(ηα+θβ+ϕγ) %